A 2250 kg car traveling to the west slows down uniformly from 22.0 m/s to 7.00 m/s. How long does it take the car to decelerate if the force on the car is 6550 N to the east?
1
Expert's answer
2020-11-02T09:19:57-0500
"F=ma=m\frac{V-v}{t}\\6550=2250\frac{22-7}{t}\\t=5.15\ s"